# Auditory Localization and Assessment of Consequential Robot Sounds: A Multi-Method Study in Virtual Reality

### v1

@@ III-A Experimental design of the localization task
Participants localized the sounds of three robots in a VR environment, providing a controlled and immersive setting that ensured consistency of auditory and visual cues. The experiment employed a within-subjects design with four independent variables: robot type (quadruped Go1, wheeled Turtlebot, wheeled HSR), AVAS (with, without AVAS), robot velocity (0.3, 0.8 m/s) and robot trajectory (head-on, radial). A total of 24 conditions (3 robots × 2 AVAS conditions × 2 velocities × 2 trajectories) was tested. Each condition was repeated 10 times per participant (240 trials total). The two trajectories were varied in blocks and their order was counterbalanced across participants. Within each block, we presented the respective combinations of robot type, AVAS and velocity in randomized order.
@@ IV-B Localization performance for radial trajectories
When the robots moved on a radial trajectory, participants localized them with an average accuracy of 10.33∘ (=7.72∘), which is comparable to head-on trials. Fig. 3a) displays that the mean absolute localization error was descriptively larger for the wheeled HSR (=12.40∘, =9.58∘) than for the quadruped Go1 (=9.39∘, =7.05∘) and the wheeled Turtlebot (=9.22∘, =6.84∘).
@@
The localization precision in radial trials was with an average intra-individual variation of 7.53∘ (=5.24∘) on a similar level as in head-on trials. The localization precision differed significantly between robots (results in Tab. IV, Fig. 3b)). The largest variation and thus the lowest precision resulted for the HSR (=8.83∘, =6.33∘). The precision for the HSR was significantly lower than both for the Go1 (=6.98∘, =5.22∘, (23)=2.68, =.020, =0.55) and the Turtlebot (=6.79∘, =4.73∘, (23)=3.70, =.004, =0.76). The localization precision did not vary significantly between Go1 and Turtlebot ((23)=0.50, =.999 =0.10).

### v2

@@ III-A Experimental design of the localization task
Participants localized the sounds of three robots in a laboratory VR environment. The use of an audiovisual VR provided a controlled and immersive setting, which allowed for precise standardization of sensory cues and improved the internal validity of the localization experiment. Specifically, it enabled us to present the exact same robot sound for a specific duration across multiple trials per participant. The experiment employed a within-subjects design with four independent variables: robot type (quadruped Go1, wheeled Turtlebot, wheeled HSR), AVAS (with, without AVAS), robot velocity (0.3, 0.8 m/s) and robot trajectory (head-on, radial). A total of 24 conditions (3 robots × 2 AVAS conditions × 2 velocities × 2 trajectories) was tested. Each condition was repeated 10 times per participant (240 trials total). The two trajectories were varied in blocks and their order was counterbalanced across participants. Within each block, we presented the respective combinations of robot type, AVAS and velocity in randomized order.
@@ IV-B Localization performance for radial trajectories
In radial trials, participants localized the robots with an average accuracy of 10.33∘ (=7.72∘), which is comparable to head-on trials. The mean absolute localization error (Fig. 3a)) was descriptively larger for the HSR (=12.40∘, =9.58∘) than for the Go1 (=9.39∘, =7.05∘) and the Turtlebot (=9.22∘, =6.84∘).
@@
The localization precision in radial trials was with an average intra-individual variation of 7.53∘ (=5.24∘) on a similar level as in head-on trials. The localization precision differed significantly between robots (results in Tab. IV, Fig. 3b)). The largest variation and thus the lowest precision resulted for the HSR (=8.83∘, =6.33∘), which was significantly lower than both for the Go1 (=6.98∘, =5.22∘, (23)=2.68, =.020, =0.55) and the Turtlebot (=6.79∘, =4.73∘, (23)=3.70, =.004, =0.76). The localization precision did not vary significantly between Go1 and Turtlebot ((23)=0.50, =.999 =0.10).